# PhysX-3D: Physical-Grounded 3D Asset Generation

### v3

@@ 5.4 Qualitative Results
Figure 5 showcases the physical-grounded 3D assets generated by our PhysXGen. By learning the interdependencies between physical and structural space, PhysXGen achieves impressive performance in generating physical properties. Besides, we perform qualitative comparisons with our baseline shown in Figure 6. For absolute scaling, material , and kinematics, we use the MAE-based similarity assessment to measure the similarity between generated and ground-truth properties. The higher score means smaller error with the ground-truth. By utilizing the interdependencies between physical properties and structural information, especially geometry, our PhysXNet obtains higher overall scores. Furthermore, our PhysXGen can distinguish the properties of different parts and achieve more stable and robust performance in physical property generation of neighboring structures, especially in function description and affordance. More experimental results are shown in the supplementary.
@@ 6 Conclusion
Appendix A Implementation details
Structure of PhysXGen. In this section, we present the architectural details and implementation specifics of PhysXGen. To maintain consistency with the established pre-trained geometry space, our geometrical decoder preserves the original hyperparameter configuration from  [27], ensuring effective utilization of pre-trained weights. For the physical processing components, we implement structurally symmetric encoder-decoder pairs (as detailed in Table 4). Notably, our physical generator employs a streamlined transformer architecture with 14 processing blocks rather than the conventional 24-block configuration to achieve satisfactory performance with lower computational overhead.
Texture retrieval for PhysXGen training. While the 3D objects in PartNet [17] inherently lack surface texturing data, we retrieve compatible UV texture coordinates from ShapeNet [3]. For instances where no corresponding texture exists in ShapeNet, we employ the grey color as their texture information.
@@ Appendix A Implementation details
Appendix B Details of human-in-the-loop annotation pipeline
@@
Methods PSNR $\uparrow$ CD $\downarrow$ F-Score $\uparrow$ Absolute scale $\downarrow$ Material $\downarrow$ Affordance $\downarrow$ Kinematic parameters $\downarrow$ Description $\downarrow$
TRE [27] + PartField [15] + GPT 24.31 13.2 76.9 8.75 0.399 0.517 1.77 0.466
PhysXGen 24.53 12.7 77.3 6.63 0.141 0.372 0.479 0.71
@@
Appendix C Procedural generation in PhysXNet-XL
@@
Appendix D More experimental results
D.1 Comparison with GPT-based baseline
To evaluate the capabilities of our proposed method, PhysXGen, in generating physically-grounded 3D assets, we conduct comprehensive qualitative and quantitative comparisons against a GPT-based baseline pipeline comprising Trellis [27], PartField [15], and GPT-4o. Under this benchmark framework, given an input image prompt, Trellis first generates textured 3D meshes with complete geometric and appearance representations. These assets are subsequently processed by PartField to perform fine-grained part segmentation, followed by a GPT-based physical property assignment module that predicts material parameters and dynamic attributes for each identified part. As shown in Table 5, our method exceeds the GPT-based method in geometry and most physics metrics. Across the four evaluation dimensions of absolute scale, material, kinematics, and affordance, PhysXGen demonstrates significant performance gains over the GPT-based baseline, achieving relative improvements of 24%, 64%, 28%, and 72%, respectively. In function description, our PhysXGen performs reduced robustness compared to GPT-4o, primarily attributable to its training on a relatively smaller dataset, i.e., PhysXNet. Furthermore, we visualize the generated results of the
@@ D.1 Comparison with GPT-based baseline
D.2 Qualitative comparison among different architectures
@@
Appendix E Further analysis on challenges in physical-grounded 3D generation
@@
Function description: Our framework leverages CLIP [20] for text embedding extraction, subsequently performing dimensionality reduction through 3D VAE to establish a compressed latent space. This architecture enables joint learning of all physical properties. However, the non-invertible nature of CLIP’s encoder-only architecture fundamentally restricts embedding-to-prompt disentanglement, thereby constraining interpretability in downstream 3D semantic reasoning tasks. Meanwhile, compared with other physical properties, text embedding is more complex to learn and generate. As shown in Fig 11, the error in normalized function descriptions is larger than other properties. Furthermore, while encoder-decoder foundation models like T5 [21] theoretically permit decoding, their high-dimensional embedding spaces impose prohibitively expensive computational overhead for cross-domain alignment with physical properties.

### v4

@@ 5.4 Qualitative Results
Figure 5 showcases the physical-grounded 3D assets generated by our PhysXGen. By learning the interdependencies between physical and structural space, PhysXGen achieves impressive performance in generating physical properties. Besides, we perform qualitative comparisons with our baseline shown in Figure 6. As we mentioned above, for absolute scaling, we use the Euclidean distance while we adopt PSNR to evaluate the material maps, affordance maps, function description similarity score maps. By utilizing the interdependencies between physical properties and structural information, especially geometry, our PhysXNet obtains higher overall scores. Furthermore, our PhysXGen can distinguish the properties of different parts and achieve more stable and robust performance in physical property generation of neighboring structures, especially in function description, material , and affordance. More experimental results are shown in the supplementary.
@@ 6 Conclusion
Acknowledgment
This study is supported by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory. This study is also supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221-0012, MOE-T2EP20223-0002). This research is also supported by cash and in-kind funding from NTU S-Lab and industry partner(s).
7 Implementation details
Structure of PhysXGen. In this section, we present the architectural details and implementation specifics of PhysXGen. To maintain consistency with the established pre-trained geometry space, our geometrical decoder preserves the original hyperparameter configuration from  trellis, ensuring effective utilization of pre-trained weights. For the physical processing components, we implement structurally symmetric encoder-decoder pairs (as detailed in Table 4). Notably, our physical generator employs a streamlined transformer architecture with 14 processing blocks rather than the conventional 24-block configuration to achieve satisfactory performance with lower computational overhead.
Texture retrieval for PhysXGen training. While the 3D objects in PartNet partnet inherently lack surface texturing data, we retrieve compatible UV texture coordinates from ShapeNet shapenet. For instances where no corresponding texture exists in ShapeNet, we employ the grey color as their texture information.
@@ 7 Implementation details
Methods Geometry Absolute scale $\downarrow$ Material $\uparrow$ Affordance $\uparrow$ Kinematic parameters Description $\uparrow$
PSNR $\uparrow$ CD $\downarrow$ F-Score $\uparrow$     COV $\uparrow$   MMD $\downarrow$
TRE trellis + PartField liu2025partfieldlearning3dfeature + GPT 24.31 13.2 76.9 8.81 7.95 6.73 0.09 0.24 14.31
PhysXGen 24.53 12.7 77.3 7.24 13.01 11.30 0.33 0.08 10.11
8 Details of human-in-the-loop annotation pipeline
@@
9 Procedural generation in PhysXNet-XL
@@
10 More experimental results
10.1 Comparison with GPT-based baseline
To evaluate the capabilities of our proposed method, PhysXGen, in generating physically-grounded 3D assets, we conduct comprehensive qualitative and quantitative comparisons against a GPT-based baseline pipeline comprising Trellis trellis, PartField liu2025partfieldlearning3dfeature, and GPT-4o. Under this benchmark framework, given an input image prompt, Trellis first generates textured 3D meshes with complete geometric and appearance representations. These assets are subsequently processed by PartField to perform fine-grained part segmentation, followed by a GPT-based physical property assignment module that predicts material parameters and dynamic attributes for each identified part. As shown in Table 5, our method exceeds the GPT-based method in geometry and most physics metrics. Across the four evaluation dimensions of absolute scale, material, kinematics, and affordance, PhysXGen demonstrates significant performance gains over the GPT-based baseline, achieving relative improvements of 24%, 64%, 28%, and 72%, respectively. In function description, our PhysXGen performs reduced robustness compared to GPT-4o, primarily attributable to its training on a relatively smaller dataset, i.e., PhysXNet. Furthermore, we visualize the generated results of the
@@ 10.1 Comparison with GPT-based baseline
10.2 Qualitative comparison among different architectures
@@
11 Further analysis on challenges in physical-grounded 3D generation
@@
Function description: Our framework leverages CLIP radford2021learning for text embedding extraction, subsequently performing dimensionality reduction through 3D VAE to establish a compressed latent space. This architecture enables joint learning of all physical properties. However, the non-invertible nature of CLIP’s encoder-only architecture fundamentally restricts embedding-to-prompt disentanglement, thereby constraining interpretability in downstream 3D semantic reasoning tasks. Meanwhile, compared with other physical properties, text embedding is more complex to learn and generate. As shown in Fig 11, the error in normalized function descriptions is larger than other properties. Furthermore, while encoder-decoder foundation models like T5 raffel2020exploring theoretically permit decoding, their high-dimensional embedding spaces impose prohibitively expensive computational overhead for cross-domain alignment with physical properties.
12 Ethics statement
Human Annotator Ethical Concerns: All annotations used in this study were performed by the authors. No external participants were involved, and no personal or sensitive data were collected. According to our institution’s ethical guidelines, this study does not constitute human subjects research and therefore did not require IRB approval.
Clarification of dataset license: Since 3D data in PhysXNet and PhysXNet-XL are derived and modified from PartNet and ShapeNet, users are required to comply with the ShapeNet license terms [^1] https://shapenet.org/terms.
Potential biases: While our dataset, PhysXNet, offers a new collection of 3D objects annotated with rich physical properties, we acknowledge that the dataset may contain representational biases, which we plan to address in future work. Moreover, we caution against the direct application of our methods in safety-critical domains (e.g., autonomous robotics or medical devices) without rigorous validation, as errors or inaccuracies in physical property predictions could result in adverse outcomes. Finally, since part of the annotation process involves vision-language models (VLMs), specifically GPT-4o (costs over \$1k), the dataset annotations may reflect biases inherent in these models despite subsequent human verification.